# USTC-ICTS/PCFT-25-44

### v2

@@ Abstract
As pointed out in recent research, the near extremal black hole entropy with one-loop effect exhibits universal $\log T$ behaviour at sufficiently low temperature. In this paper, we discuss the low temperature quantum corrections to the thermodynamics of four dimensional accelerating black holes with rotation and charges by using the method of Euclidean path integral. The one-loop path integral for the black hole thermal partition function near extremality is dominated by zero modes defined with respect to the extremal background. For the accelerating black holes without rotation, the near horizon extremal geometry is a direct product of AdS2 and S2 with warping factors, and the gravitational zero modes contain both tensor and vector types, with the respective contributions to the near extremal black hole entropy being $(3/2)\log T$ and $(1/2)\log T$. While in the presence of rotation, the near horizon extremal geometry is a twist product of AdS2 and S2 and the gravitational vector modes are absent. For the accelerating black holes with charges, we also consider the one-loop path integral of the gauge field, where the photon zero modes are found to contribute an additional $(1/2)\log T$ term to the near extremal black hole entropy.
@@ 1 Introduction
Black holes are fascinating thermal objects with their own laws of thermodynamics that are analogous to those of classical thermodynamics. From the semiclassical analysis, the temperature is determined by the surface gravity and the entropy is proportional to the area of the horizon of the black hole Bekenstein:1973ur ; Hawking:1976de ; Gibbons:1976ue ; York:1986it . However, black holes are distinct from ordinary thermal systems. As two of the Killing horizons converge, the temperature of a charged or rotating black hole can become very small, while the entropy remains finite value since the black hole horizon has non-vanishing area in the extremal limit. This phenomenon indicates there exists a large amount of ground state degeneracy with the density of black hole states proportional to the exponential of its horizon area. For supersymmetric black holes in string theory, the thermal entropy in the extremal cases are well understood by the microscopic state counting Strominger:1996sh ; Sen:2014aja . For non-supersymmetric black holes like those in our universe, such a large degeneracy need further explanations. Whether there exists an unknown symmetry protecting mechanism at low temperature to ensure the ground state degeneracy or it is simply a result of the semiclassical approximation remains an open question for thorough investigation.
@@
One step forward is to incorporate quantum corrections into the calculation of the black hole entropy when going beyond the semiclassical approximation in the gravitational path integral. For extremal black holes, the leading quantum correction to the entropy is found to take a universal form of logarithm of the horizon area Banerjee:2010qc ; Banerjee:2011jp ; Sen:2012kpz ; Sen:2012cj , without invoking the explicit ultraviolet structure of the underlying quantum theory of gravity. In the non-extremal cases, a complete profile of the finite temperature quantum correction still hide in mist. Once the temperature is very small, the near horizon geometry of any near extremal black hole in four dimensions universally takes the form as a product of AdS2 and S2. With this property, the quantum fluctuations above the extremal or near extremal configurations exhibit controllable behaviours. As pointed out in Iliesiu:2020qvm ; Heydeman:2020hhw ; Boruch:2022tno ; Iliesiu:2022onk ; Rakic:2023vhv , the quantum fluctuations in the near horizon region of higher dimensional extremal or near extremal black holes are well captured by the two dimensional Jackiw-Teitelboim (JT) gravity Jackiw:1984je ; Teitelboim:1983ux through dimensional reduction and the low energy physics is described by a Schwarzian theory Almheiri:2014cka ; Maldacena:2016upp with certain boundary conditions of the AdS2 factor in the near horizon geometry. By using the path integral of the Schwarzian theory at low temperature, $\log T$ corrections are found in the thermal entropy of the theory, which is different from the logarithm of the horizon area correction to the extremal entropy. The $\log T$ term in entropy is essential to resolve the mass gap puzzle Preskill:1991tb appears in the low temperature region of non-extremal black holes under the semiclassical approximation. Without quantum corrections, the black hole thermal energy above extremality scales quadratically in small temperature while the average energy of the Hawking radiation scales linearly in temperature. This means that the black hole can not radiate even a single Hawking quantum at sufficiently low temperature. The $\log T$ quantum correction to the entropy in turn corrects the thermal energy above extremality with a term linear in small temperature, and thus address the issue involved in the mass gap puzzle.
@@
The $\log T$ corrections can also be figured out by using the Euclidean path integral without resorting to the dimensional reduction. In reference Banerjee:2023quv , the authors evaluate the one-loop partition function for the near extremal black hole in the Einstein-Maxwell theory by using the Euclidean path integral. A set of zero modes with zero eigenvalues of the kinetic operator of small fluctuations are found to contribute leading quantum corrections in the path integral. In their analysis, the near extremal black holes are viewed as small temperature deviations form the extremal ones. The small temperature perturbation regularize the divergent contribution from the zero modes path integral on the extremal background and result a $\log T$ term in the logarithm of the partition function. The same logic also applies to the Kerr black holes Kapec:2023ruw , where the near horizon extremal Kerr (NHEK) geometry plays a crucial role in determining the zero modes in the Euclidean path integral. Interestingly, these zero modes are generated by the vector fields which are asymptotic to those large diffeomorphisms in the Kerr/CFT correspondence Guica:2008mu ; Castro:2009jf ; Bredberg:2011hp . See Maulik:2024dwq ; Maulik:2025phe for the discussions of the universality of the low temperature quantum corrections for the asymptotically AdS and dS black holes by using the gravitational path integral method. For charged black holes with additional gauge fields, the quantum corrections from the gauge fields path integral should also be taken into account. It is found in  Blacker:2025zca that the gauge zero modes also contribute in the Euclidean path integral and lead to additional $\log T$ corrections to the thermal entropy. Thus, the coefficient of the $\log T$ term depends on the near horizon geometry of the extremal black hole as well as the field contents in the gravity theory.
@@
In this paper, we will follow the spirit of Kapec:2023ruw and consider the low temperature quantum corrections for accelerating black holes. An accelerating black hole is usually known as the C-metric which describes a pair of black holes constantly accelerating away from each other Kinnersley:1970zw ; Plebanski:1976gy ; Dias:2002mi ; Hong:2003gx ; Hong:2004dm ; Griffiths:2005se ; Griffiths:2005qp ; Podolsky:2006px . Without specifying the range of the polar angle, the metric of an accelerating black hole contains deficit angles at the north and south poles. Therefore, its acceleration can be effectively viewed as the result of cosmic strings attached at the poles with tensions determined by the deficit angles. Much like the Kerr case, a rotating black hole with acceleration also has an infinite near horizon region at extremality. For example, the extremal accelerating Kerr-Newmann black hole is shown to acquire a warped and twisted product of AdS2 and S2 near horizon scaling region, where the asymptotic symmetry analysis can be applied to realize a 2D CFT dual Astorino:2016xiy . The thermodynamic properties of accelerating black holes have been widely explored Appels:2016uha ; Astorino:2016ybm ; Appels:2017xoe ; Anabalon:2018ydc ; Zhang:2018hms ; Anabalon:2018qfv ; Gregory:2019dtq ; Ball:2020vzo ; Ball:2021xwt ; Cassani:2021dwa ; Kim:2023ncn ; Wu:2023meo ; Liu:2025iyl ; Hale:2025veb . We will use the Euclidean path integral to figure out the low temperature quantum corrections to the thermodynamics of accelerating black holes, considering both rotating and charged cases. The existences of the AdS2 factor in their near horizon extremal geometry entails that a similar statement regarding the low temperature quantum corrections holds true in these cases. We will highlight the effect due to the acceleration in comparing with the non-accelerating cases.
@@
This paper is organized as follows. In section 2, we first review the the thermodynamics of the accelerating Kerr black hole, and the low temperature behaviours of the thermodynamic quantities. We carry out the near horizon geometry of the near extremal accelerating Kerr black hole with an infinite scaling coordinate transformation and solve the zero modes of the kinetic operator in the one-loop Euclidean path integral on the extremal background. Then we perturb the extremal background with the first order temperature perturbation and calculate the finite temperature corrections to the eigenvalue of the zero modes. These finite and non-vanishing eigenvalues determine the leading one-loop quantum corrections to the logarithm of the partition function as $(3/2)\log T$. Section 3 is for the case where there is a gauge field present while the black hole remains static, i.e., the accelerating Reissner-Nordström black hole. In the absence of rotation, the near horizon extremal geometry is a direct product of AdS2 and S2 with warping factors, and the gravitational zero modes correspondingly are divided into tensor and vector modes. We introduce the finite temperature perturbation to the eigenvalue problem of the tensor and vector modes and find their contributions to the one-loop path integral of the logarithm of the partition function as $(3/2)\log T$ and $(1/2)\log T$, respectively. In addition, the gauge field fluctuation contains photon zero modes and their path integral gives another $(1/2)\log T$ term to the logarithm of the partition function. In section 4, we consider the accelerating Kerr-Newmann black hole. The presence of rotation sets the near horizon extremal geometry as a warped and twisted product of AdS2 and S2. Similar to the accelerating Kerr case, the gravitational zero modes defined on the extremal background only contains tensor type modes, and their contribution to the logarithm of the partition function is $(3/2)\log T$. The gauge field fluctuation also contains photon zero modes and produce a $(1/2)\log T$ term to the logarithm of the partition function. With all the three cases analyzed, we can see a pattern that the zero modes structure is highly dependent on the near horizon extremal geometry as well as the field contents in the theory. The $\log T$ quantum correction to the logarithm of the partition function at low temperature is a universal result with coefficient $3/2$ resulting from each tensor type zero modes and $1/2$ from each vector type zero modes.
@@ 2 Accelerating Kerr black hole
We begin with the accelerating Kerr black hole, namely, a four dimensional rotating black hole with acceleration. This is a solution to the pure Einstein gravity, with its metric in Boyer-Lindquist coordinates taking the following form Hong:2004dm ; Griffiths:2005se ; Griffiths:2005qp
@@
where $\Theta_{\pm}=1\mp 2\alpha m+\alpha^{2}a^{2}$ and we set the period of the polar angle $\phi$ to be $2\pi$. The deficit angles $\delta_{\pm}=2\pi-2\pi\Theta_{\pm}$ at the poles are effectively induced by the cosmic strings attached on the poles with tensions $\mu_{\pm}=\delta_{\pm}/(8\pi)$. In terms of the black hole parameters, we have
@@ 2.1 Near extremal thermodynamics and geometry
The accelerating Kerr black hole (2.1) is parameterized by $m$, $a$, $\alpha$, so there are three independent thermodynamic quantities that can be used to describe all the other thermodynamic quantities. We take this triple as the temperature $T$, angular momentum $J$ and net string tension $\mu$. The black hole temperature $T$ in (2.7) vanishes in the extremal case. A small temperature therefore parameterizes a small deviation from extremality. The angular momentum $J_{0}=r_{0}^{2}$ in (2.7) and the net string tension $\mu_{0}=r_{0}\alpha$ (2.5) are freely chosen as fixed at their extremal values. Thus in the near extremal case, the other thermodynamic quantities can be expanded as power series in terms of the small temperature. By using relations (2.6) and (2.7), we find that the horizon radius $r_{+}$ and $r_{-}$ can be expanded as
@@
Note that the outer horizon is inside the accelerating horizon, so we have $0<\mu_{0}=r_{0}\alpha<1$. The first term on the right side of (2.12) is understood as the the thermal energy at extremality, while the second term is the leading thermal energy above extremality which scales quadratically with the temperature of the black hole. However, the typical energy of a quantum of Hawking radiation scales linearly with the temperature. Below the temperature where these two energy scales meet, the semiclassical analysis gives problematic low temperature behaviours of the thermodynamic quantities since the emission of a single Hawking quantum could carry away all of the thermal energy available in the black hole system. This is known as the mass gap puzzle first discussed in Preskill:1991tb .
In order to resolve this puzzle, one need to go beyond the semiclassical analysis and introduce vital quantum corrections to the thermodynamics at low temperature. From the perspective of thermal partition function, quantum corrections are introduced by the Euclidean path integral over quantum fluctuations around a given semiclassical saddle-point. Since we are considering the quantum corrections for low temperature thermodynamics, the saddle-point here should be chosen as the extremal geometry of the black hole. To obtain the extremal geometry, we perform the following scaling transformation
@@
This transformation enables us to zoom into the infinite near horizon region of the near extremal accelerating Kerr black hole with a small temperature. Using the expansions (2.10), (2.11) and the coordinate transformation (2.14), the black hole metric (2.1) can be written as a power series of $T$. The leading $\mathcal{O}(T^{0})$ term of the metric describe the extremal geometry, which is
@@
The metric (2.16) describes the near horizon extremal geometry of the accelerating Kerr black hole, which is a warped and twist product of AdS2 and S2. The near horizon near extremal geometry $d\bar{s}^{2}+\delta(ds^{2})$ is obtained by adding the first order low temperature perturbation (2.17). In the following discussion, the near extremal configuration will be used as the perturbed background in the Euclidean path integral and determines the leading quantum corrections to the thermodynamic quantities of the accelerating Kerr black hole.

### v3

@@ Abstract
As pointed out in recent research, the near extremal black hole entropy with one-loop effect exhibits universal $\log T$ behavior at sufficiently low temperature. In this paper, we discuss the low-temperature quantum corrections to the thermodynamics of four-dimensional accelerating black holes with rotation and charges by using the method of Euclidean path integral. The one-loop path integral for the black hole thermal partition function near extremality is dominated by zero modes defined with respect to the extremal background. For the accelerating black holes without rotation, the near horizon extremal geometry is a direct product of AdS2 and S2 with warping factors, and the gravitational zero modes contain both tensor and vector types, with the respective contributions to the near extremal black hole entropy being $(3/2)\log T$ and $(1/2)\log T$. While in the presence of rotation, the near horizon extremal geometry is a warped and twisted product of AdS2 and S2 and the gravitational vector modes are absent. For the accelerating black holes with charges, we also consider the one-loop path integral of the gauge field, where the photon zero modes are found to contribute an additional $(1/2)\log T$ term to the near extremal black hole entropy.
@@ 1 Introduction
Black holes are fascinating thermal objects with their own laws of thermodynamics that are analogous to those of classical thermodynamics. From the semiclassical analysis, the temperature is determined by the surface gravity and the entropy is proportional to the area of the horizon of the black hole Bekenstein:1973ur ; Hawking:1976de ; Gibbons:1976ue ; York:1986it . However, black holes are distinct from ordinary thermal systems. As two of the Killing horizons converge, the temperature of a charged or rotating black hole can become very small, while the entropy remains finite value since the black hole horizon has nonvanishing area in the extremal limit. This phenomenon indicates there exists a large amount of ground-state degeneracy with the density of black hole states proportional to the exponential of its horizon area. For supersymmetric black holes in string theory, the thermal entropy in the extremal cases are well-understood by the microscopic state counting Strominger:1996sh ; Sen:2014aja . For nonsupersymmetric black holes like those in our Universe, such a large degeneracy need further explanations. Whether there exists an unknown symmetry protecting mechanism at low temperature to ensure the ground state degeneracy or it is simply a result of the semiclassical approximation remains an open question for thorough investigation.
@@
One step forward is to incorporate quantum corrections into the calculation of the black hole entropy when going beyond the semiclassical approximation in the gravitational path integral. For extremal black holes, the leading quantum correction to the entropy is found to take a universal form of logarithm of the horizon area Banerjee:2010qc ; Banerjee:2011jp ; Sen:2012kpz ; Sen:2012cj , without invoking the explicit ultraviolet structure of the underlying quantum theory of gravity. In the nonextremal cases, a complete profile of the finite temperature quantum correction still hide in mist. Once the temperature is very small, the near horizon geometry of any near extremal black hole in four dimensions universally takes the form as a product of two dimensional anti-de Sitter space (AdS2) and two sphere (S2). With this property, the quantum fluctuations above the extremal or near extremal configurations exhibit controllable behaviors. As pointed out in Iliesiu:2020qvm ; Heydeman:2020hhw ; Boruch:2022tno ; Iliesiu:2022onk ; Rakic:2023vhv , the quantum fluctuations in the near horizon region of higher-dimensional extremal or near extremal black holes are well captured by the two-dimensional Jackiw-Teitelboim (JT) gravity Jackiw:1984je ; Teitelboim:1983ux through dimensional reduction and the low energy physics is described by a Schwarzian theory Almheiri:2014cka ; Maldacena:2016upp with certain boundary conditions of the AdS2 factor in the near horizon geometry. By using the path integral of the Schwarzian theory at low temperature, $\log T$ corrections are found in the thermal entropy of the theory, which is different from the logarithm of the horizon area correction to the extremal entropy. The $\log T$ term in entropy is essential to resolve the mass gap puzzle Preskill:1991tb appears in the low-temperature region of nonextremal black holes under the semiclassical approximation. Without quantum corrections, the black hole thermal energy above extremality scales quadratically in small temperature while the average energy of the Hawking radiation scales linearly in temperature. This means that the black hole can not radiate even a single Hawking quantum at sufficiently low temperature. The $\log T$ quantum correction to the entropy in turn corrects the thermal energy above extremality with a term linear in small temperature, and thus address the issue involved in the mass gap puzzle.
@@
The $\log T$ corrections can also be figured out by using the Euclidean path integral without resorting to the dimensional reduction. In Ref. Banerjee:2023quv , the authors evaluate the one-loop partition function for the near extremal black hole in the Einstein-Maxwell theory by using the Euclidean path integral. A set of zero modes with zero eigenvalues of the kinetic operator of small fluctuations are found to contribute leading quantum corrections in the path integral. In their analysis, the near extremal black holes are viewed as small temperature deviations form the extremal ones. The small temperature perturbation regularize the divergent contribution from the zero modes path integral on the extremal background and result a $\log T$ term in the logarithm of the partition function. The same logic also applies to the Kerr black holes Kapec:2023ruw , where the near horizon extremal Kerr (NHEK) geometry plays a crucial role in determining the zero modes in the Euclidean path integral. Interestingly, these zero modes are generated by the vector fields which are asymptotic to those large diffeomorphisms in the Kerr/CFT correspondence Guica:2008mu ; Castro:2009jf ; Bredberg:2011hp . See Maulik:2024dwq ; Maulik:2025phe for the discussions of the universality of the low-temperature quantum corrections for the asymptotically AdS and dS black holes by using the gravitational path integral method. For charged black holes with additional gauge fields, the quantum corrections from the gauge fields path integral should also be taken into account. It is found in  Blacker:2025zca that the gauge zero modes also contribute in the Euclidean path integral and lead to additional $\log T$ corrections to the thermal entropy. Thus, the coefficient of the $\log T$ term depends on the near horizon geometry of the extremal black hole as well as the field contents in the gravity theory.
@@
In this paper, we will follow the spirit of Kapec:2023ruw and consider the low-temperature quantum corrections for accelerating black holes. An accelerating black hole is usually known as the C-metric which describes a pair of black holes constantly accelerating away from each other Kinnersley:1970zw ; Plebanski:1976gy ; Dias:2002mi ; Hong:2003gx ; Hong:2004dm ; Griffiths:2005se ; Griffiths:2005qp ; Podolsky:2006px . Without specifying the range of the polar angle, the metric of an accelerating black hole contains deficit angles at the north and south poles. Therefore, its acceleration can be effectively viewed as the result of cosmic strings attached at the poles with tensions determined by the deficit angles. Much like the Kerr case, a rotating black hole with acceleration also has an infinite near horizon region at extremality. For example, the extremal accelerating Kerr-Newmann black hole is shown to acquire a warped and twisted product of AdS2 and S2 near horizon scaling region, where the asymptotic symmetry analysis can be applied to realize a 2D conformal field theory (CFT) dual Astorino:2016xiy . The thermodynamic properties of accelerating black holes have been widely explored Appels:2016uha ; Astorino:2016ybm ; Appels:2017xoe ; Anabalon:2018ydc ; Zhang:2018hms ; Anabalon:2018qfv ; Gregory:2019dtq ; Ball:2020vzo ; Ball:2021xwt ; Cassani:2021dwa ; Kim:2023ncn ; Wu:2023meo ; Liu:2025iyl ; Hale:2025veb . We will use the Euclidean path integral to figure out the low-temperature quantum corrections to the thermodynamics of accelerating black holes, considering both rotating and charged cases. The existences of the AdS2 factor in their near horizon extremal geometry entails that a similar statement regarding the low-temperature quantum corrections holds true in these cases. We will highlight the effect due to the acceleration in comparing with the nonaccelerating cases.
@@
This paper is organized as follows. In Sec. 2, we first review the the thermodynamics of the accelerating Kerr black hole, and the low-temperature behaviors of the thermodynamic quantities. We carry out the near horizon geometry of the near extremal accelerating Kerr black hole with an infinite scaling coordinate transformation and solve the zero modes of the kinetic operator in the one-loop Euclidean path integral on the extremal background. Then we perturb the extremal background with the first-order temperature perturbation and calculate the finite temperature corrections to the eigenvalue of the zero modes. These finite and nonvanishing eigenvalues determine the leading one-loop quantum corrections to the logarithm of the partition function as $(3/2)\log T$. Section 3 is for the case where there is a gauge field present while the black hole remains static, i.e., the accelerating Reissner-Nordström black hole. In the absence of rotation, the near horizon extremal geometry is a direct product of AdS2 and S2 with warping factors, and the gravitational zero modes correspondingly are divided into tensor and vector modes. We introduce the finite temperature perturbation to the eigenvalue problem of the tensor and vector modes and find their contributions to the one-loop path integral of the logarithm of the partition function as $(3/2)\log T$ and $(1/2)\log T$, respectively. In addition, the gauge field fluctuation contains photon zero modes and their path integral gives another $(1/2)\log T$ term to the logarithm of the partition function. In Sec. 4, we consider the accelerating Kerr-Newmann black hole. The presence of rotation sets the near horizon extremal geometry as a warped and twisted product of AdS2 and S2. Similar to the accelerating Kerr case, the gravitational zero modes defined on the extremal background only contains tensor type modes, and their contribution to the logarithm of the partition function is $(3/2)\log T$. The gauge field fluctuation also contains photon zero modes and produce a $(1/2)\log T$ term to the logarithm of the partition function. With all the three cases analyzed, we can see a pattern that the zero modes structure is highly dependent on the near horizon extremal geometry as well as the field contents in the theory. The $\log T$ quantum correction to the logarithm of the partition function at low temperature is a universal result with coefficient $3/2$ resulting from each tensor type zero modes and $1/2$ from each vector type zero modes.
@@ 2 Accelerating Kerr black hole
We begin with the accelerating Kerr black hole, namely, a four-dimensional rotating black hole with acceleration. This is a solution to the pure Einstein gravity, with its metric in Boyer-Lindquist coordinates taking the following form Hong:2004dm ; Griffiths:2005se ; Griffiths:2005qp
@@
where $\Theta_{\pm}=1\mp 2\alpha m+\alpha^{2}a^{2}$ and we set the period of the polar angle $\phi$ to be $2\pi$. Note that we can introduce a constant factor $K$ to the polar angle $\phi/K$ so that one of the conical singularities can be removed. As we will see, such a constant will be canceled in the calculation of the eigenvalues of the zero modes since the zero modes are normalized. So, here we keep the range of the polar angle as $2\pi$ for simplicity. The deficit angles $\delta_{\pm}=2\pi-2\pi\Theta_{\pm}$ at the poles are effectively induced by the cosmic strings attached on the poles with tensions $\mu_{\pm}=\delta_{\pm}/(8\pi)$. In terms of the black hole parameters, we have
@@ 2.1 Near extremal thermodynamics and geometry
The accelerating Kerr black hole (2.1) is parametrized by $m$, $a$, $\alpha$, so there are three independent thermodynamic quantities that can be used to describe all the other thermodynamic quantities. We take this triple as the temperature $T$, angular momentum $J$ and net string tension $\mu$. The black hole temperature $T$ in (2.7) vanishes in the extremal case. A small temperature therefore parametrizes a small deviation from extremality. The angular momentum $J_{0}=r_{0}^{2}$ in (2.7) and the net string tension $\mu_{0}=r_{0}\alpha$ (2.5) are freely chosen as fixed at their extremal values. Thus in the near extremal case, the other thermodynamic quantities can be expanded as power series in terms of the small temperature. By using relations (2.6) and (2.7), we find that the horizon radius $r_{+}$ and $r_{-}$ can be expanded as
@@
Note that the outer horizon is inside the accelerating horizon, so we have $0<\mu_{0}=r_{0}\alpha<1$. The first term on the right side of (2.12) is understood as the the thermal energy at extremality, while the second term is the leading thermal energy above extremality which scales quadratically with the temperature of the black hole. However, the typical energy of a quantum of Hawking radiation scales linearly with the temperature. Below the temperature where these two energy scales meet, the semiclassical analysis gives problematic low-temperature behaviors of the thermodynamic quantities since the emission of a single Hawking quantum could carry away all of the thermal energy available in the black hole system. This is known as the mass gap puzzle first discussed in Preskill:1991tb .
In order to resolve this puzzle, one need to go beyond the semiclassical analysis and introduce vital quantum corrections to the thermodynamics at low temperature. From the perspective of thermal partition function, quantum corrections are introduced by the Euclidean path integral over quantum fluctuations around a given semiclassical saddle-point. Since we are considering the quantum corrections for low-temperature thermodynamics, the saddle-point here should be chosen as the extremal geometry of the black hole. To obtain the extremal geometry, we perform the following scaling transformation
@@
This transformation enables us to zoom into the infinite near horizon region of the near extremal accelerating Kerr black hole with a small temperature. Using the expansions (2.10), (2.11), and the coordinate transformation (2.14), the black hole metric (2.1) can be written as a power series of $T$. The leading $\mathcal{O}(T^{0})$ term of the metric describe the extremal geometry, which is
@@
The metric (2.16) describes the near horizon extremal geometry of the accelerating Kerr black hole, which is a warped and twisted product of AdS2 and S2. The near horizon near extremal geometry $d\bar{s}^{2}+\delta(ds^{2})$ is obtained by adding the first-order low-temperature perturbation (2.17). In the following discussion, the near extremal configuration will be used as the perturbed background in the Euclidean path integral and determines the leading quantum corrections to the thermodynamic quantities of the accelerating Kerr black hole.